a 1490 kg car has just passed a diner. the closest the car came to the diner was r=40.5, but the car has driven another x=30.7 m in the 2.25 s elapsed since then. If the car travels at constant speed, what is the current magnitude of the angular momentum of the car about the diner?
Answer #1

30.27 m in 2.25 s

so speed of car = 30.27/2.25 = 13.45 m/s

so angular momentum = r*m*v= 40.5*1490*13.45 = 811640.35
Answer #2

v=30.7/2.25=13.644 m/s


==> L=mvR=13.644*1490*40.5=823374 kg-m^2/s
Answer #3

823,373.73